Bukit Barisan (07.01.2010 to 09.30.2010)

You are here

PDF versionPDF version
Year: 
2010
Quarter: 
July - September
Communications activities: 

No news

Unusual Events at the TEAM Site: 

No news

New Species: 

No news

Protocol Activities: 

Climate: All system are running well, but we still have problems in retrieving the data every two weeks because of limited staffs who is mobile and able to use internet to send the data. Vegetation: All vegetation plots (VGBBS 1, 2, 3, 4, 5 and 6) have been set up and all tree/liana have been measured during May-August 2010. In each plot 477, 436, 376, 480, 223 an 450 stem were measured respectively. Camera Trap: Our second camera trap (CT-BBS-02) has been set-up and pick-up. All the data in process of identifying species.

Protocol Problems: 

No news

Schedule Problems: 

No news

Logistical Problems: 

No logistical problem at this time